A block of mass M is pulled with a rope that exert
Answer #1

[1] ans

the force on X-axis=>Fx=Fcos\theta

the force on y-axis =>Fy=F sin\theta

[2] ans

the force on y-axis=>Fy=mg-Fsin\theta+N

[3] ans

the normal force N=mg-Fsin\theta

[4] ans

the force on X-axis=>Fx=\mu[mg-Fsin\theta)

[5] ans

acceleration a=\mug/[cos\theta+\musin\theta)